OpenStudy (anonymous):

Your Math class has 32 students. You would like to treat the class to cookies. You would like each person to have 5 cookies. Which expression would be a quick way to figure out how many cookies you would need to bake? 5(32 + 2) 5(30 + 2) 5(32 − 2) 5(30 − 2)

OpenStudy (anonymous):

Well, we know 5 time 32 would be our answer so what expression you have that will give you 5 * 32 ???

OpenStudy (anonymous):

Hint: 5 cookies per person. and there are 32 students. 5x32..:)

OpenStudy (anonymous):

out of my choices above bc thats what i did 5x32

OpenStudy (anonymous):

Ok you have 5 * 32. That is the answer but what does 5 times 32 equal?

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

YW

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

I made a boo boo on my last post. It was meant as the following 5(30 − 2) 5(30) = 150 5(-2) = -10 150 -10 = 140

OpenStudy (anonymous):

so thats my answer

OpenStudy (anonymous):

No, you are looking for an expression that equals 160 because 5 time 32 = 160 5(30 − 2) = 140 so this one is not it Try the other ones out and see which one equals 160.

OpenStudy (anonymous):

You try and tell me what you get and I will tell you if your are right or wrong

OpenStudy (anonymous):

You know 5(30 − 2) is not it. I did that one for you

OpenStudy (anonymous):

alright :)

OpenStudy (anonymous):

Why you are so confused? You just need the answer which gives 32 inside brackets.. which is B..:P

OpenStudy (anonymous):

the answer is 5(30 + 2) bc it equals 160! :)

OpenStudy (anonymous):

Yes ashlee good job

OpenStudy (anonymous):

yay!